# Failure of the measure contraction property via quotients in higher-step sub-Riemannian structures

### v1

@@ 8. Failure of the measure contraction property for the Martinet structure
Since the Martinet structure admits dilations, it is enough to disprove $\mathrm{MCP}(0,N)$. We show that we get a contradiction by setting $\omega=2\operatorname{K}(m)$ in (8.5) and sending $m\to 1$. Using the special values of the Jacobi elliptic functions [AS64, 16.8] we obtain for $m\in(0,1)$,
$\mathcal{J}_{R}(2\operatorname{K}(m),m)=-4\left[\operatorname{E}(m)^{2}+(m-1)\operatorname{K}(m)\big{(}2\operatorname{E}(m)-\operatorname{K}(m)\big{)}\right].$ (243)
Furthermore with the help of [AS64, 16.16] we get
$\partial_{\omega}\mathcal{J}_{R}(2\operatorname{K}(m),m)=2\Big{[}(1-2m)\operatorname{E}(m)+(m-1)\operatorname{K}(m)\Big{]}.$ (244)
Using [AS64, 17.3.26 and 17.3.27] we have
$\lim_{m\to 1}\operatorname{K}(m)=+\infty,\quad\lim_{m\to 1}\operatorname{E}(m)=1,\quad\lim_{m\to 1}(m-1)\operatorname{K}(m)=\lim_{m\to 1}(m-1)\operatorname{K}(m)^{2}=0.$ (245)
Therefore, we obtain
$\lim_{m\to 1}1+2\operatorname{K}(m)\frac{\partial_{\omega}\mathcal{J}_{R}(2\operatorname{K}(m),m)}{\mathcal{J}_{R}(2\operatorname{K}(m),m)}=1+2(+\infty)\frac{(-2)}{(-4)}=+\infty,$ (246)
implying by Proposition 8.5 that the Martinet structure does not satisfy the $\mathrm{MCP}(0,N)$. ∎
@@ A.1. The Goh-Legendre condition
We refer to [ABB20, Ch. 12] for details, we only note here that all abnormal geodesics that are not normal must have finite index so that they must be Goh-Legendre abnormals.
We also formulate a strengthening of the Goh-Legendre condition. It was introduced in [AS96, (4.16)] in relation with rigidity properties of abnormal geodesics. We emphasize that the (strong) generalized Legendre condition should not be confused with the (strong) Legendre condition in control theory, see [AS04, Ch. 20 and in particular Ch. 20.5.1].
@@ Appendix B Approximations of probability measures
$\left|\int_{X}g(z)\mu^{m}(\mathrm{d}z)-\int_{X}g(z)\mu(\mathrm{d}z)\right|$ $=\left|\sum_{i=1}^{L(m)}\mu\left(A_{i}^{m}\right)g(y_{i}^{m})-\int_{X}g(z)\mu(\mathrm{d}z)\right|$ (269)

### v2

@@ 9. Failure of the $\mathrm{MCP}$ for the Martinet structure
Since the Martinet structure admits dilations, it is enough to disprove $\mathrm{MCP}(0,N)$. We show that we get a contradiction in (240) by writing an expansion of $\mathcal{J}_{R}$ around $m=1$ and letting $\omega$ tend to +$\infty$+ ∞. The following expansion can be found in [AS64, 16.15.1–3]:
$\operatorname{sn}(\xi\mid m)$ $=\tanh(\xi)+\frac{1}{4}(m-1)\left(\sinh(\xi)\cosh(\xi)-\xi\right)\operatorname{sech}(\xi)^{2}+O\big{(}(m-1)^{2}\big{)};$ (242)
$\operatorname{cn}(\xi\mid m)$ $=\operatorname{sech}(\xi)-\frac{1}{4}(m-1)(\sinh(\xi)\cosh(\xi)-\xi)\tanh(\xi)\operatorname{sech}(\xi)+O\big{(}(m-1)^{2}\big{)};$ (243)
$\operatorname{dn}(\xi\mid m)$ $=\operatorname{sech}(\xi)+\frac{1}{4}(m-1)(\sinh(\xi)\cosh(\xi)+\xi)\tanh(\xi)\operatorname{sech}(\xi)+O\big{(}(m-1)^{2}\big{)}.$ (244)
Recalling from [AS64, 17.2.10] that $\operatorname{E}(\xi\mid m):=\int_{0}^{\xi}\operatorname{dn}(u\mid m)^{2}\mathrm{d}u$, we also obtain
$\operatorname{E}(\xi\mid m)=\tanh(\xi)-\frac{1}{4}(m-1)\Big{[}\xi+\tanh(\xi)(\xi\tanh(\xi)-1)\Big{]}+O\big{(}(m-1)^{2}\big{)}.$
Plugging these formulas into (234) gives
$\mathcal{J}_{R}(\xi,m):=\frac{1}{2}(m-1)\cosh(\xi)\Big{[}\sinh(\xi)^{2}+\xi\left(\tanh(\xi)-2\xi\right)\Big{]}+O\big{(}(m-1)^{2}\big{)}.$
Therefore, we get, for all $t\in(0,1)$, that
$\lim_{\omega\to+\infty}\lim_{m\to 1^{-}}\left\lvert\frac{\mathcal{J}_{R}(\omega t,m)}{\mathcal{J}_{R}(\omega,m)}\right\rvert=\lim_{\omega\to\infty}\frac{\cosh(\omega t)}{\cosh(\omega)}\frac{\sinh^{2}(\omega t)+\omega t\left(\tanh(\omega t)-2\omega t\right)}{\sinh^{2}(\omega)+\omega\left(\tanh(\omega)-2\omega\right)}=0.$ (245)
This implies by Proposition 9.5 that the Martinet structure does not satisfy the $\mathrm{MCP}(0,N)$. ∎
Remark 9.7 (Violation of weaker conditions).
Suppose that there is a function $f:[0,1]\to[0,+\infty)$ such that for all $q\in\mathbb{Mar}$ and all Borel set $A\subseteq\mathbb{Mar}$, with $0<\mathsf{m}(A)<\infty$ it holds
$\mathsf{m}(Z_{t}(q,A))\geq f(t)\mathsf{m}(A),\qquad\forall\,t\in[0,1].$ (246)
The same argument leading to (241) yields that for all $t\in[0,1]$, it holds
$t|\mathcal{J}_{R}(\omega t,m)|\geq f(t)|\mathcal{J}_{R}(\omega,m)|,\qquad\forall\,m\in(0,1)\,,\ \forall\,\omega\in(0,2\operatorname{K}(m)).$
The computation (245) shows that $f(t)\equiv 0$ for all $t\in[0,1)$. The Martinet structure is therefore not qualitatively non-degenerate in the sense of [CH15, Assumption 1].
10. Variants and related curvature conditions
In this paper, we work with Ohta’s $\mathrm{MCP}$, the weakest among the variants of the measure contraction property. A variant in terms of the Rényi entropy was used by Cavalletti and Mondino in [CM17] (see also [CM21, Def. 6.8], where it is denoted by $\mathrm{MCP}_{\varepsilon}$). One can deduce that $\mathrm{MCP}_{\varepsilon}$ implies Ohta’s $\mathrm{MCP}$, see [CM21, Lemma 6.13]. Since we do not assume the essentially non-branching condition, the converse implication might not be true.
Furthermore, for $Q\in[1,\infty)$, one can define a relaxed version of the $\mathrm{MCP}$ (resp. $\mathrm{MCP}_{\varepsilon}$) by multiplying the inequality (14) in Definition 2.1 (resp. [CM21, (6.4) in Def. 6.8]) by a factor $1/Q$. The case $Q=1$ corresponds to Ohta’s $\mathrm{MCP}$ and the $\mathrm{MCP}_{\varepsilon}$ variant.
With minor modifications in the proofs, all the results concerning stability under quotients that we proved for $\mathrm{MCP}$ also hold for the $Q$-relaxed variants and the corresponding $\mathrm{MCP}_{\varepsilon}$. Arguing without loss of generality in the case $K=0$, any such variant easily implies that
$\mathsf{m}(Z_{t}(x,A))\geq\frac{t^{N}}{Q}\mathsf{m}(A),\qquad\forall\,t\in[0,1],$ (247)
for the $t$-intermediate set between $x\in X$ and a Borel set $A$ with $0<\mathsf{m}(A)<+\infty$. Condition (247) cannot be satisfied by the Martinet structure; see Remark 9.7.
We conclude that any sub-Riemannian metric measure space which, as a consequence of Theorem 1.9, does not satisfy the $\mathrm{MCP}$, also fails to satisfy the $\mathrm{MCP}_{\varepsilon}$ and the corresponding $Q$-relaxed variants. This includes, of course, all red Carnot groups appearing in Table 1.
Finally, one can verify with an approximation argument that the Quasi Curvature-Dimension condition $\mathrm{QCD}(Q,K,N)$ introduced in [Mil21], for $Q\in[1,\infty)$, implies the $Q$-relaxed version of the $\mathrm{MCP}_{\varepsilon}(K,N)$ mentioned above. Thus, again under the conditions of Theorem 1.9, Milman’s $\mathrm{QCD}$ condition cannot be satisfied either.
@@ A.1. The Goh-Legendre condition
We refer to [ABB20, Ch. 12] for details; we only note here that all abnormal geodesics that are not normal must have finite index so that they must be Goh-Legendre geodesics.
We also formulate a strengthening of the Goh-Legendre condition. It was introduced in [AS96, (4.16)] in relation with rigidity properties of abnormal geodesics.
@@ Appendix B Approximations of probability measures
$\Big{|}\int_{X}g(z)\mu^{m}(\mathrm{d}z)$ $-\int_{X}g(z)\mu(\mathrm{d}z)\Big{|}=\left|\sum_{i=1}^{L(m)}\mu\left(A_{i}^{m}\right)g(y_{i}^{m})-\int_{X}g(z)\mu(\mathrm{d}z)\right|$ (270)